OpenStudy (anonymous):

Simplify, and remember to rationalize the denominator if needed. Show your work as much as you can for partial credit. √16a divided by √4a

OpenStudy (mertsj):

\[\frac{\sqrt{16a}}{\sqrt{4a}}=\sqrt{\frac{16a}{4a}}\]

OpenStudy (mertsj):

Perhaps you can finish it.